(x+3)squared=8(y+2)
Is this ellipse, parabola or hyperbola

Answer by venugopalramana(3286)

(x+3)squared=8(y+2)
Is this ellipse, parabola or hyperbola
PARABOLA...
TIP IS IF THERE IS ONLY X^2 OR Y^2 TERM IT COULD BE PARABOLA.
STD.EQN.OF PARABOLA IS
(X-H)^2=4A(Y-K) WHICH FITS INTO THE ABOVE EQN.